9. Find the center of gravity of a system consisting of two spheres of masses 500g and 800g, welded to the ends of the rod of negligible mass with length 60cm. The radii of the spheres are 10cm and 15cm respectively.


1
Expert's answer
2021-03-14T19:15:52-0400

"\sum \tau=0\to m_1\cdot x=m_2\cdot(l-x)"


"l=0.1+0.6+0.15=0.85\ (m)"


"0.5\cdot x=0.8\cdot(0.85-x)\to x=0.523\ (m)"


"0.523 \ (m)" from the center of a sphere of mass "500\ (g)" . Answer